# Kernel-based Koopman approximants for control: Flexible sampling, error analysis, and stability

### v1

@@ 2.2. Approximation error for regularized kEDMD: uniform bounds
In this section, we extend recently proposed uniform error bounds for the approximant $\widehat{\mathcal{K}}$ in the operator norm $\|\cdot\|_{\mathbb{H}\to C_{b}(\Omega)}$ to regularized kEDMD. Hereby, we focus on Wendland kernels generating the RKHS $\mathcal{N}_{\Phi_{n,k}}(\Omega)$ (cf. Example 2.2), analogously to [20], and refer to Remark 2.6 for a discussion on a potential alternative.
If the size of the data set $\mathcal{X}$ is large, the evaluation of Equation 7 may lead to numerical instabilities as the kernel matrix $\mathbf{K}_{\mathcal{X}}$ is typically badly conditioned. For this reason, one often regularizes kernel-based interpolation problems of the form Equation 5. More precisely, the regression problem is endowed by a regularization term, i.e.,
$R^{\lambda}_{\mathcal{X}}f:=\operatorname*{arg\,min}_{g\in V_{\mathcal{X}}}\;\sum_{i=1}^{d}|f(x_{i})-g(x_{i})|^{2}+\lambda\|g\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)}^{2},$
where $\lambda\geq 0$ is the regularization parameter. Clearly, $R_{\mathcal{X}}^{0}=P_{\mathcal{X}}$ and it can be easily verified that the solution operator $R_{\mathcal{X}}^{\lambda}$, $\lambda\geq 0$, is linear and satisfies
$R^{\lambda}_{\mathcal{X}}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}},\qquad f\in\mathcal{N}_{\Phi_{n,k}}(\Omega).$
However, while $P_{\mathcal{X}}$ is a projection, the operator $R^{\lambda}_{\mathcal{X}}$, $\lambda>0$, is not. Nevertheless, as it is proved in Proposition B.1, $R_{\mathcal{X}}^{\lambda}$, $\lambda\geq 0$, shares with $P_{\mathcal{X}}$ that it is self-adjoint and positive semi-definite as an operator on $\mathcal{N}_{\Phi_{n,k}}(\Omega)$. Moreover, $R_{\mathcal{X}}^{\lambda}$, $\lambda\geq 0$, commutes with $P_{\mathcal{X}}$. For regularized kEDMD, we define the following approximant of the Koopman operator:
@@
where $\lambda\geq 0$. It is not hard to see that $\widehat{\mathcal{K}}_{\lambda}=R_{\mathcal{X}}^{\lambda}\mathcal{K}$. Moreover, the approximation Equation 6 is recovered for $\lambda=0$, that is, $\widehat{\mathcal{K}}_{0}=\widehat{\mathcal{K}}$
The next theorem provides a novel bound on the approximation error. This result extends the previous work [20] which considered the case $\lambda=0$, i.e., non-regularized interpolation. For a data set $\mathcal{X}\subset\Omega$ as in Equation 1, we denote the fill distance by
$h_{\mathcal{X}}:=\sup_{x\in\Omega}\,\operatorname{dist}(x,\mathcal{X})$
[…]
@@
$|(\mathcal{K}f)(x)-(\widehat{\mathcal{K}}_{\lambda}f)(x)|\leq C\big{(}h_{\mathcal{X}}^{k+1/2}+\sqrt{\lambda}\big{)}\|f\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)}.$
We require an auxiliary result based on [47, Proposition 3.6] to prove Theorem 2.4.
@@
Let $k\in\mathbb{N}$. Then there are constants $C,h_{0}>0$ such that for every finite set $\mathcal{X}=\{x_{j}\}_{j=1}^{d}\subset\Omega$ of sample points with $h_{\mathcal{X}}\leq h_{0}$ and all multiindices $\alpha\in\mathbb{N}_{0}^{n}$, $|\alpha|\leq k-\frac{(-1)^{n}+1}{2}$, we have for all $\lambda\geq 0$, $f\in\mathcal{N}_{\Phi_{n,k}}(\Omega)$ and $x\in\Omega$
@@
The result [47, Proposition 3.6] states that if $\tau=\ell+s$ with $\ell\in\mathbb{N}$, $\ell>n/2$, $s\in(0,1]$, and $20\leq j<\ell-n/$ 2, then for $f\in H^{\tau}(\Omega)$ we have
$|f-R_{\mathcal{X}}^{\lambda}f|_{W^{j,\infty}(\Omega)}\,\leq\,C\Big{(}h_{\mathcal{X}}^{\tau-j-n/2}+h_{\mathcal{X}}^{-j}\sqrt{\lambda}\Big{)}\|f\|_{H^{\tau}(\Omega)},$
where $|\cdot|_{W^{j,\infty}(\Omega)}$ is the semi-norm $|u|_{W^{j,\infty}(\Omega)}=\sup_{|\alpha|=j}\|D^{\alpha}u\|_{L^{\infty}(\Omega)}$.
Let us choose $\tau=\sigma_{n,k}=\frac{n+1}{2}+k$. If $n$ is odd, we have to set $\ell=\tau$ and $s=0$. Then $\ell-n/2=k+1/2$ so that, if $|\alpha|\leq k$, setting $j=|\alpha|$ is possible. Hence, $\tau-j-n/2=k+1/2-|\alpha|$, and the desired estimate follows. On the other hand, if $n$ is even, then $\ell=k+n/2$ and $s=1/2$ so that only $10\leq j\leq k-$ 1 is allowed. Hence, if $|\alpha|\leq k-1$, the result follows. ∎
@@
The proof of the theorem follows the lines of that of [20, Theorem 5.2] together with Lemma 2.5. ∎
@@
It is well known that the native space of the Matérn kernel
$\mathsf{k}(x,y)$ $=\frac{2^{1-\nu-n/2}}{\pi^{\nu/2}\Gamma(\nu)\sigma^{2\nu-n}}\big{(}\sigma\|x-y\|\big{)}^{\nu-n/2}K_{\nu-n/2}\big{(}\sigma\|x-y\|\big{)},\qquad x,y\in\mathbb{R}^{n}$
is the Sobolev space $H^{\nu}(\mathbb{R}^{n})$, cf. [7, Example 5.7]. Here, $\Gamma$ is the gamma function, $K_{q}$ the modified Bessel function of the second kind, and $\sigma>0$, $\nu\in\mathbb{N}$, $\nu>n/2$. Since (under suitable regularity properties of the flow) Sobolev spaces are invariant under the Koopman operator, a similar result as Theorem 2.4 holds for kEDMD.
@@ 2.3. Data-driven surrogate dynamics
In what follows, we propose a Koopman-based surrogate model for the dynamics Equation DS. For this, let $\psi_{j}\in\mathcal{N}_{\Phi_{n,k}}(\Omega)$, $j\in[1:M]$, be observables such that the map $\Psi=[\psi_{1},\ldots,\psi_{M}]^{\top}\in\mathcal{N}_{\Phi_{n,k}}(\Omega)^{M}$ has a left inverse $\Upsilon:\mathbb{R}^{M}\to\mathbb{R}^{n}$, i.e., $\Upsilon(\Psi(x))=x$ for $x\in\Omega$, with modulus of continuity $\omega_{\Upsilon}\in\mathscr{K}$. Note that continuity and injectivity of $\Psi$ imply $M\geq n$. This follows, for example, from the Borsuk-Ulam theorem, see, e.g., [42, Theorem V.8.9].
@@
Since the right-hand side of the dynamics Equation DS satisfies
@@
$x^{+}=\widehat{F}_{\lambda}(x)=\Upsilon\big{(}\big{[}\widehat{\mathcal{K}}_{\lambda}\psi_{1}(x),\ldots,\widehat{\mathcal{K}}_{\lambda}\psi_{M}(x)\big{]}^{\top}\big{)}=\Upsilon\big{(}\Psi_{F(\mathcal{X})}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}(x)\big{)}$ (10)
with $\Psi_{\mathcal{Z}}=[\Psi(z_{1}),\ldots,\Psi(z_{d})]^{\top}\in\mathbb{R}^{d\times M}$, where we utilize the approximation $\widehat{K}_{\lambda}$, $\lambda\geq 0$, of the Koopman operator provided in Equation 8. We note that if the flow map $F$ can be evaluated directly, one may choose $\Psi=\Upsilon=\operatorname{id}$.
@@
Making use of Theorem 2.4, we compute
@@
$=\omega_{\Upsilon}\bigg{(}\bigg{[}\sum_{j=1}^{M}\big{|}\mathcal{K}\psi_{j}(x)-\widehat{\mathcal{K}}_{\lambda}\psi_{j}(x)\big{|}^{2}\bigg{]}^{1/2}\bigg{)}$
$\leq\omega_{\Upsilon}\bigg{(}C\left(h_{\mathcal{X}}^{k+1/2}+\sqrt{\lambda}\right)\bigg{[}\sum_{j=1}^{M}\|\psi_{j}\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)}^{2}\bigg{]}^{1/2}\bigg{)},$
which yields the claim. ∎
@@ 3. Data-driven surrogates: Lyapunov stability
In this part, we provide our first main result. We provide sufficient conditions that ensure transferability of stability results from the original dynamics to the data-driven surrogate and vice versa. To be more precise, we show that asymptotic stability of the dynamical system Equation DS, certified by a Lyapunov function $V$ with modulus of continuity $\omega_{V}$, implies semi-global practical asymptotic stability of the kEDMD surrogate model given by Equation 10 and vice versa. Furthermore, under some compatibility assumptions on the Lyapunov function, semi-global asymptotic stability is even preserved for the unregularized surrogate, i.e., $\widehat{F}_{\lambda}$ with $\lambda=0$.
To keep the presentation technically simple(r), in the following we consider the unregularized data-driven surrogate dynamics
@@
as defined in Equation 10 and provide the details for the regularized surrogate dynamics, i.e., $\widehat{F}_{\lambda}$ with $\lambda>0$, in subsequent comments. Due to the relation $\widehat{\mathcal{K}}_{0}=\widehat{\mathcal{K}}$, this corresponds to vanilla kernel EDMD Equation 6.
An important ingredient for the subsequent stability analysis is that equilibria are preserved in the data-driven surrogate models if they are contained in the set $\mathcal{X}$ of data points. We recall that a state $x^{*}\in\Omega$ is called an equilibrium of the dynamics Equation DS if $F(x^{*})=x^{*}$.
@@
A data point $x^{*}\in\mathcal{X}$ is an equilibrium of the dynamics Equation DS if and only if it is an equilibrium of the surrogate dynamics Equation 12.
@@
For each $k\in[1:d]$, we have
$\Psi_{F(\mathcal{X})}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x_{k})=\Psi_{F(\mathcal{X})}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X}}e_{k}=\Psi_{F(\mathcal{X})}^{\top}e_{k}=\Psi(F(x_{k}))$
Applying $\Upsilon$, the left inverse of $\Psi$, to both sides of this equation yields $\widehat{F}(x_{k})=F(x_{k})$ for all $k\in[1:d]$. Thus, as $x^{*}\in\mathcal{X}$, we have $F(x^{*})=\widehat{F}(x^{*})$ which implies the claim. ∎
We point out that the relation of Proposition 3.1 above is only approximately preserved when considering regularized kEDMD Equation 10 with a regularization parameter $\lambda>0$. More precisely, in view of the implicit function theorem, the data-driven surrogate has an equilibrium in a neighborhood of the original model’s equilibrium and the size of this neighborhood is proportional to ${\lambda}$.
@@
Consider the discrete-time dynamical system Equation DS on the set $\Omega$. Let $Y\subset\Omega$ be a forward invariant set w.r.t. the dynamics Equation DS containing the equilibrium $x^{*}=F(x^{*})$ in its interior $\operatorname{int}(Y)$. Then, we have the following definitions:
@@
(ii) Let, in addition, a set $P\subset Y$ be given that is forward invariant w.r.t. the dynamics Equation DS. The origin is called $P$-practically asymptotically stable on $Y$ if there exists $\beta\in\mathscr{K\!L}$ such that, for all $x\in Y$ and $n\in\mathbb{N}_{0}$, either the inclusion $F^{n}(x)\in P$ or inequality Equation 13 holds.
Asymptotic stability can be characterized by means of Lyapunov functions, see, e.g., [11] and the references therein.
@@
A continuous function $V:Y\subset\Omega\to\mathbb{R}_{\geq 0}$ is said to be a Lyapunov function w.r.t. the dynamics Equation DS and the equilibrium $x^{*}=F(x^{*})\in Y$ if there exist $\alpha_{1},\alpha_{2}\in\mathscr{K}_{\infty}$ and $\alpha_{V}\in\mathscr{K}$ such that
@@
$V(F(x))\leq V(x)-\alpha_{V}(\|x-x^{*}\|)$ (15)
holds for all $x\in Y$ with $F(x)\in Y$.
The following proposition is assembled from [10, Theorem 2.19 and 2.20].
@@
Let the sets $P,Y\subset\Omega$ satisfying $P\subset Y$ be forward invariant w.r.t. the dynamics Equation DS with equilibrium $x^{*}\in\operatorname{int}(P)$. If $V$ is a Lyapunov function in accordance to Definition 3.3, the equilibrium $x^{*}$ is asymptotically stable on $Y$. Alternatively, if $V$ is a Lyapunov function satisfying the decrease condition Equation 15 on $S=Y\backslash P$, $x^{*}$ is $P$-practically asymptotically stable on $Y$.

### v3

@@ 2.2. Approximation error for regularized kEDMD: uniform bounds
In this section, we extend recently proposed uniform error bounds for the approximant $\widehat{\mathcal{K}}$ in the operator norm $\|\cdot\|_{\mathbb{H}\to C_{b}(\Omega)}$ to regularized kEDMD.
If the size of the data set $\mathcal{X}$ is large, the evaluation of eq. 6 may lead to numerical instabilities as the kernel matrix $\mathbf{K}_{\mathcal{X}}$ is typically badly conditioned. For this reason, one often regularizes kernel-based interpolation problems of the form eq. 4. More precisely, the regression problem is endowed by a regularization term, i.e.,
$R^{\lambda}_{\mathcal{X}}f:=\operatorname*{arg\,min}\nolimits_{g\in V_{\mathcal{X}}}\;\sum\nolimits_{i=1}^{d}|f(x_{i})-g(x_{i})|^{2}+\lambda\|g\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)}^{2}$
with regularization parameter $\lambda\geq 0$. Clearly, $R_{\mathcal{X}}^{0}=P_{\mathcal{X}}$ holds and it can be verified that the solution operator $R_{\mathcal{X}}^{\lambda}$ is linear and satisfies $R^{\lambda}_{\mathcal{X}}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}$, $f\in\mathcal{N}_{\Phi_{n,k}}(\Omega)$. However, while $P_{\mathcal{X}}$ is a projection, the operator $R^{\lambda}_{\mathcal{X}}$, $\lambda>0$, is not. Nevertheless, as proved in the (additional) appendix (see LABEL:{p:RX}), $R_{\mathcal{X}}^{\lambda}$, $\lambda\geq 0$, shares with $P_{\mathcal{X}}$ that it is self-adjoint and positive semi-definite as an operator on $\mathcal{N}_{\Phi_{n,k}}(\Omega)$. Moreover, $R_{\mathcal{X}}^{\lambda}$, $\lambda\geq 0$, commutes with $P_{\mathcal{X}}$. For regularized kEDMD, we define the following approximant of the Koopman operator:
@@
where $\lambda\geq 0$. It is not hard to see that $\widehat{\mathcal{K}}_{\lambda}=R_{\mathcal{X}}^{\lambda}\mathcal{K}$. Moreover, the approximation eq. 5 is recovered for $\lambda=0$, that is, $\widehat{\mathcal{K}}_{0}=\widehat{\mathcal{K}}$.
The next theorem provides a novel bound on the approximation error. This result extends the previous work [28], which considered the case $\lambda=0$, i.e., non-regularized interpolation. For a data set $\mathcal{X}\subset\Omega$ as in eq. 1, we denote the fill distance by
$h_{\mathcal{X}}:=\sup\nolimits_{x\in\Omega}\,\operatorname{dist}(x,\mathcal{X})$
using the Euclidean norm $\|\cdot\|$ as the metric for the distance operator. Analogously to [28], and refer to remark 2.6 for a discussion on a potential alternative.
@@
Let $k\geq 1$, $\lambda\geq 0$ and $F\in C_{b}^{\lceil\sigma_{n,k}\rceil}(\Omega;\mathbb{R}^{n})$. Then there are constants $C,h_{0}>0$ such that for any finite set $\mathcal{X}=\{x_{i}\}_{i=1}^{d}\subset\Omega$ of pairwise distinct sample points with $h_{\mathcal{X}}\leq h_{0}$ and for all $f\in\mathcal{N}_{\Phi_{n,k}}(\Omega)$ and $x\in\Omega$, we have
$|(\mathcal{K}f)(x)-(\widehat{\mathcal{K}}_{\lambda}f)(x)|\leq C\big(h_{\mathcal{X}}^{k+1/2}+\sqrt{\lambda}\big)\|f\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)},$
where the RKHS $\mathcal{N}_{\Phi_{n,k}}(\Omega)$ (cf. example 2.2) is generated by Wendland kernels.
We require an auxiliary result based on [56, Proposition 3.6] to prove theorem 2.4.
@@
Let $k\in\mathbb{N}$. Then there are constants $C,h_{0}>0$ such that for every finite set $\mathcal{X}=\{x_{j}\}_{j=1}^{d}\subset\Omega$ of sample points with $h_{\mathcal{X}}\leq h_{0}$ and all multiindices $\alpha\in\mathbb{N}_{0}^{n}$, $|\alpha|\leq k-1$, we have for all $\lambda\geq 0$, $f\in\mathcal{N}_{\Phi_{n,k}}(\Omega)$ and $x\in\Omega$
@@
[56, Proposition 3.6] states that if $\tau=\ell+s$ with $\ell\in\mathbb{N}$, $\ell>n/2$, $s\in(0,1]$, and $0\leq j<\ell-n/2$, then $|f-R_{\mathcal{X}}^{\lambda}f|_{W^{j,\infty}(\Omega)}\,\leq\,C\Big(h_{\mathcal{X}}^{\tau-j-n/2}+h_{\mathcal{X}}^{-j}\sqrt{\lambda}\Big)\|f\|_{H^{\tau}(\Omega)}$ holds for $f\in H^{\tau}(\Omega)$, where $|\cdot|_{W^{j,\infty}(\Omega)}$ is the semi-norm $|u|_{W^{j,\infty}(\Omega)}=\sup_{|\alpha|=j}\|D^{\alpha}u\|_{L^{\infty}(\Omega)}$.
Let us choose $\tau=\sigma_{n,k}=\frac{n+1}{2}+k$. If $n$ is odd, we have to set $\ell=\tau-1$ and $s=1$. Then $\ell-n/2=k-1/2>0$ so that, if $|\alpha|\leq k-1$, setting $j=|\alpha|$ is possible. Hence, $\tau-j-n/2=k+1/2-|\alpha|$, and the desired estimate follows. On the other hand, if $n$ is even, then $\ell=k+n/2$ and $s=1/2$ so that only $0\leq j\leq k-1$ is allowed. Hence, if $|\alpha|\leq k-1$, the result follows. ∎
@@
The proof of the theorem follows the lines of that of [28, Theorem 5.2] together with lemma 2.5. ∎
Let us briefly discuss the convergence of $\widehat{\mathcal{K}}_{0}$ to $\mathcal{K}$ in the infinite-data limit, that is, $h_{\mathcal{X}}\to 0$ (implying $|\mathcal{X}|\to\infty$). theorem 2.4 establishes convergence in the operator norm $\|\cdot\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)\to C_{b}(\Omega)}$ and provides quantitative bounds. However, this result does not imply convergence in operator norms, such as $\|\cdot\|_{C_{b}(\Omega)\to C_{b}(\Omega)}$ or $\|\cdot\|_{L^{2}(\Omega)\to L^{2}(\Omega)}$, nor does it ensure any form of spectral convergence. For detailed discussions on spectral convergence as well as more sophisticated EDMD variants, see, e.g., [8] and [9].
@@
The proof of theorem 2.4 on the Koopman approximation consists of two steps. The first ingredient is Koopman invariance of the RKHS, that is, $\mathcal{K}\mathbb{H}\subset\mathbb{H}$. Note that this property has to be checked carefully for chosen kernels. While for Wendland kernels corresponding to fractional Sobolev spaces (and thus also for Matérn kernels generating the same spaces [12, Example 5.7]), this invariance holds as shown in [28], there are also negative results, e.g., for Gaussian kernels [13]. The second step is establishing an interpolation (or finite-data) error estimate as stated in lemma 2.5, e.g., for general radial basis functions in [36, 37].
@@ 2.3. Data-driven surrogate dynamics
In what follows, we propose a Koopman-based surrogate model for the dynamics eq. DS. For this, let $\psi_{j}\in\mathcal{N}_{\Phi_{n,k}}(\Omega)$, $j\in[1:M]$, be observables such that the map $\Psi=[\psi_{1},\ldots,\psi_{M}]^{\top}\in\mathcal{N}_{\Phi_{n,k}}(\Omega)^{M}$ has a left inverse $\Upsilon:\mathbb{R}^{M}\to\mathbb{R}^{n}$, i.e., $\Upsilon(\Psi(x))=x$ for $x\in\Omega$, with modulus of continuity $\omega_{\Upsilon}\in\mathscr{K}$. Note that continuity and injectivity of $\Psi$ imply $M\geq n$. This follows, for example, from the Borsuk-Ulam theorem, see, e.g., [51, Theorem V.8.9]. The left inverse map $\Upsilon$ can be chosen linear if and only if the coordinate projections $x\mapsto x_{i}$, $i\in[1:n]$, are in the linear span of the lifting $\Psi$.
@@
Since the right-hand side of the dynamics eq. DS satisfies
@@
$x^{+}=\widehat{F}_{\lambda}(x)=\Upsilon\big(\big[\widehat{\mathcal{K}}_{\lambda}\psi_{1}(x),\ldots,\widehat{\mathcal{K}}_{\lambda}\psi_{M}(x)\big]^{\hskip-2.84526pt\top}\big)=\Upsilon\big(\Psi_{F(\mathcal{X})}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}(x)\big)$ (9)
with $\Psi_{\mathcal{Z}}=[\Psi(z_{1}),\ldots,\Psi(z_{d})]^{\top}\in\mathbb{R}^{d\times M}$, where we utilize the approximation $\widehat{K}_{\lambda}$, $\lambda\geq 0$, of the Koopman operator provided in eq. 7. We note that if the flow map $F$ can be evaluated directly, one may choose $\Psi=\Upsilon=\operatorname{id}$.
@@
Making use of theorem 2.4, the claim follows from the computation
@@
$=\omega_{\Upsilon}\big(\big[\sum\nolimits_{j=1}^{M}\big|\mathcal{K}\psi_{j}(x)-\widehat{\mathcal{K}}_{\lambda}\psi_{j}(x)\big|^{2}\big]^{1/2}\big)$
$\leq\omega_{\Upsilon}\big(C\left(h_{\mathcal{X}}^{k+1/2}+\sqrt{\lambda}\right)\big[\sum\nolimits_{j=1}^{M}\|\psi_{j}\|_{\mathcal{N}_{\Phi_{n,k}}(\Omega)}^{2}\big]^{1/2}\big).$
@@ 3. Data-driven surrogates: Lyapunov stability
In this part, we provide our first main result. We provide sufficient conditions that ensure transferability of stability results from the original dynamics to the data-driven surrogate and vice versa. To be more precise, we show that asymptotic stability of the dynamical system eq. DS, certified by a Lyapunov function $V$ with modulus of continuity $\omega_{V}$, implies semi-global practical asymptotic stability of the kEDMD surrogate model given by eq. 9 and vice versa. Furthermore, under some compatibility assumptions on the Lyapunov function, semi-global asymptotic stability is even preserved for the unregularized surrogate, i.e., $\widehat{F}_{\lambda}$ with $\lambda=0$. To keep the presentation technically simple(r), in the following we consider the unregularized data-driven surrogate dynamics
@@
and discuss the regularized surrogate dynamics ($\widehat{F}_{\lambda}$ with $\lambda>0$), in subsequent comments. Due to the relation $\widehat{\mathcal{K}}_{0}=\widehat{\mathcal{K}}$, this corresponds to vanilla kernel EDMD eq. 5.
An important ingredient for the subsequent stability analysis is that equilibria are preserved in the surrogate models if they are contained in the set $\mathcal{X}$ of data points. We recall that a state $x^{*}\in\Omega$ is called an equilibrium of the dynamics eq. DS if $F(x^{*})=x^{*}$.
@@
A data point $x^{*}\in\mathcal{X}$ is an equilibrium of the dynamics eq. DS if and only if it is an equilibrium of the surrogate dynamics eq. 11.
@@
For each $k\in[1:d]$, we have $\Psi_{F(\mathcal{X})}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x_{k})=\Psi_{F(\mathcal{X})}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X}}e_{k}=\Psi_{F(\mathcal{X})}^{\top}e_{k}=\Psi(F(x_{k})),$ where $e_{k}$ denotes the $k$th canonical unit vector. Applying $\Upsilon$, the left inverse of $\Psi$, to both sides of this equation yields $\widehat{F}(x_{k})=F(x_{k})$ for all $k\in[1:d]$. Thus, as $x^{*}\in\mathcal{X}$, we have $F(x^{*})=\widehat{F}(x^{*})$ which implies the claim. ∎
We point out that the relation of proposition 3.1 above is only approximately preserved when considering regularized kEDMD eq. 9 with a regularization parameter $\lambda>0$. More precisely, in view of the implicit function theorem, the data-driven surrogate has an equilibrium in a neighborhood of the original model’s equilibrium and the size of this neighborhood is proportional to $\lambda$}.
@@
Consider the discrete-time dynamical system eq. DS on the set $\Omega$. Let $x^{*}$ be an equilibrium, i.e., $x^{*}=F(x^{*})$, and let $Y\subset\Omega$ be a forward invariant set w.r.t. the dynamics eq. DS containing the equilibrium $x^{*}$ in its interior $\operatorname{int}(Y)$. Then,
@@
(ii) Let, in addition, a set $P\subset Y$ be given that is forward invariant w.r.t. the dynamics eq. DS. The equilibrium $x^{*}$ is called $P$-practically asymptotically stable on $Y$ if there exists $\beta\in\mathscr{K\!L}$ such that, for all $x\in Y$ and $n\in\mathbb{N}_{0}$, either the inclusion $F^{n}(x)\in P$ or inequality eq. 12 holds.
If the set $P$ is a ball centered at the equilibrium $x^{*}$ with radius $r$, practical asymptotic stability may simply be defined by
$\|F^{n}(x)-F^{n}(x^{*})\|\leq\max\{r,\beta(\|x-x^{*}\|,n)\}\quad\text{for all $x\in Y$ and $n\in\mathbb{N}_{0}$.}$ (13)
Asymptotic stability can be characterized by means of Lyapunov functions, see [16].
@@
A continuous function $V:Y\subset\Omega\to\mathbb{R}_{\geq 0}$ is said to be a Lyapunov function w.r.t. the dynamics eq. DS and the equilibrium $x^{*}=F(x^{*})\in Y$ if there exist $\alpha_{1},\alpha_{2}\in\mathscr{K}_{\infty}$ and $\alpha_{V}\in\mathscr{K}$ such that
@@
$V(F(x))\leq V(x)-\alpha_{V}(\|x-x^{*}\|)\qquad\forall\,x\in Y\text{ with }F(x)\in Y.$ (15)
The following proposition is assembled from [15, Theorem 2.19 and 2.20].
@@
Let the sets $P,Y\subset\Omega$ satisfying $P\subset Y$ be forward invariant w.r.t. the dynamics eq. DS with equilibrium $x^{*}\in\operatorname{int}(P)$. Then, the following holds:
  1. (i)
    If $V$ is a Lyapunov function in accordance to definition 3.3, the equilibrium $x^{*}$ is asymptotically stable on $Y$.
  2. (ii)
[…]